# Q-balls Under Spontaneously Broken U(1) Symmetry

### v1

@@ Abstract
We study the evolution of Q-balls under a spontaneously broken global $U(1)$ symmetry. Q-balls are stabilized by the conservation of $U(1)$ charge, but when the symmetry is spontaneously broken, the resulting Nambu–Goldstone (NG) boson can carry charge away from the Q-ball, potentially leading to charge leakage. To study this process in a controlled setting, we consider a scenario where Q-balls first form under an unbroken $U(1)$ symmetry, which is then spontaneously broken. We introduce two complex scalar fields: one responsible for forming the Q-ball, and the other for spontaneously breaking the $U(1)$ symmetry, allowing us to clearly separate the formation and symmetry-breaking phases. Using numerical simulations in a spherically symmetric system, we find that the evolution of Q-balls depends sensitively on the structure of the interaction between the two fields and the magnitude of symmetry breaking. Depending on parameters, Q-balls can completely decay, evaporate into smaller, stable Q-balls, or transition into oscillons/I-balls. In particular, we find that stable, localized remnants often survive the evolution over long timescales, especially when the symmetry-breaking scale is small. These results demonstrate that, even though spontaneous $U(1)$ breaking can lead to significant energy and charge loss from Q-balls, stable localized objects with reduced or no charge can frequently survive and potentially contribute to cosmological relics.
@@ I Introduction
Baryon and lepton numbers are typical examples of such approximate global $U(1)$ symmetries. They are well conserved at low energies, and can support the formation and stability of Q-balls. However, they are expected to be violated at high energies—for instance, through non-renormalizable operators. In the Affleck-Dine (AD) mechanism [3, 4], one of the flat directions in the minimal supersymmetric standard model (MSSM) carries baryon and/or lepton number and receives an explicit $U(1)$-breaking A-term in the potential. This A-term gives a kick in the phase direction, initiating rotation in field space and generating a net baryon and/or lepton charge. Depending on the potential shape, the flat direction may experience spatial instabilities to fragment into Q-balls [5, 6, 7, 8, 9, 10, 11]. Since the $U(1)$ breaking A-term becomes more effective when the field amplitude is large, its impact is particularly significant when Q-balls carry large charge or grow in charge after formation. In such cases, the A-term could destabilize Q-balls or hinder the growth of the Q-ball charge [12, 13, 14, 15]. The (in)stability of Q-balls affects, for instance, the viability of Q-ball dark matter [6, 16, 17] and the constraint on Q-balls by the observations of neutron stars [18, 19]. Thus, the stability of Q-balls depends critically on how well the underlying global $U(1)$ symmetry is preserved.
@@ III.3 $(n_{1},n_{2})=(3,1)$
We show the field configuration and the field trajectories at $r=0$ for $\eta=0.5$ in Fig. 6. While $\Phi$ concentrates near the center, its shape is perturbed compared with the initial Gaussian configuration, and the trajectory significantly deviates from a circle. $\Psi$ is also perturbed near the center with smaller amplitudes.
@@ IV Summary
The purpose of this work is to show the variety of the evolution of a Q-ball in a system with the spontaneously broken $U(1)$ and to establish a basis for further research on their cosmological implications. However, we briefly mention possible cosmological implications. For instance, when Q-balls decay or deform to smaller Q-balls or oscillons, their charge and energy are radiated to the universe. While these radiations can modify the thermal history of the universe or affect cosmological observables such as big bang nucleosynthesis, cosmic microwave background, and large-scale structure, the remaining objects can serve as dark matter. In addition, gravitational waves can be emitted at that time although it requires deviation from spherical symmetry in the evolution of Q-balls. Further investigation in these directions is left for future studies.

### v2

@@ Abstract
We study the evolution of Q-balls under a spontaneously broken global $U(1)$ symmetry. Q-balls are stabilized by the conservation of $U(1)$ charge, but when the symmetry is spontaneously broken, the resulting Nambu-Goldstone (NG) boson can carry charge away from the Q-ball, potentially leading to charge leakage. To study this process in a controlled setting, we consider a scenario where Q-balls first form under an unbroken $U(1)$ symmetry, which is then spontaneously broken. We introduce two complex scalar fields: one responsible for forming the Q-ball, and the other for spontaneously breaking the $U(1)$ symmetry, allowing us to clearly separate the formation and symmetry-breaking phases. Using numerical simulations in a spherically symmetric system, we find that the evolution of Q-balls depends sensitively on the structure of the interaction between the two fields and the magnitude of symmetry breaking. Depending on parameters, Q-balls can completely decay, evaporate into smaller, stable Q-balls, or transition into oscillons/I-balls. In particular, we find that stable, localized remnants often survive the evolution over long timescales, especially when the symmetry-breaking scale is small. These results demonstrate that, even though spontaneous $U(1)$ breaking can lead to significant energy and charge loss from Q-balls, stable localized objects with reduced or no charge can frequently survive and potentially contribute to cosmological relics.
@@ I Introduction
Baryon and lepton numbers are typical examples of such approximate global $U(1)$ symmetries. They are well conserved at low energies, and can support the formation and stability of Q-balls. However, they are expected to be violated at high energies—for instance, through non-renormalizable operators. In the Affleck-Dine (AD) mechanism [3, 4], one of the flat directions in the minimal supersymmetric standard model (MSSM) carries baryon and/or lepton number and receives an explicit $U(1)$-breaking A-term in the potential. This A-term gives a kick in the phase direction, initiating rotation in field space and generating a net baryon and/or lepton charge. Depending on the potential shape, the flat direction may experience spatial instabilities to fragment into Q-balls [5, 6, 7, 8, 9, 10, 11, 12]. Since the $U(1)$ breaking A-term becomes more effective when the field amplitude is large, its impact is particularly significant when Q-balls carry large charge or grow in charge after formation. In such cases, the A-term could destabilize Q-balls or hinder the growth of the Q-ball charge [13, 14, 15, 16]. The (in)stability of Q-balls affects, for instance, the viability of Q-ball dark matter [6, 17, 18] and the constraint on Q-balls by the observations of neutron stars [19, 20]. Thus, the stability of Q-balls depends critically on how well the underlying global $U(1)$ symmetry is preserved.
@@ III.3 $(n_{1},n_{2})=(3,1)$
[…]
@@ IV Summary
The purpose of this work is to show the variety of the evolution of a Q-ball in a system with the spontaneously broken $U(1)$ and to establish a basis for further research on their cosmological implications. However, we briefly mention possible cosmological implications. In our analysis, we have assumed Q-balls as initial conditions without any explicit or spontaneous breaking of the associated global $U(1)$ symmetry. Nevertheless, our results can be applied to Q-balls formed via the AD mechanism where the $U(1)$ symmetry is explicitly broken by the so-called A-terms. This is because A-terms typically originate from higher-dimensional operators suppressed by a high-energy scale, and their effects become negligibly small once the field amplitude decreases in the expanding universe. By the time Q-balls form, the effect of the A-term has typically become negligible, as it affects the dynamics only near the onset of oscillations of the corresponding flat direction. Furthermore, even in the absence of $U(1)$ charge asymmetry, the oscillating scalar field is known to evolve into Q-balls with both positive and negative charges [38, 39]. One can also consider a scenario involving multiple flat directions, where Q-balls are formed from the oscillation of one flat direction, and at a later stage, another flat direction develops a nonzero VEV. This leads to spontaneous $U(1)$ symmetry breaking, which can give rise to the types of effects discussed in this work. Thus, we can discuss the cosmological implications of our findings for a broad range of Q-ball formation mechanisms. For instance, when Q-balls decay or deform to smaller Q-balls or oscillons, their charge and energy are radiated to the universe. While these radiations can modify the thermal history of the universe or affect cosmological observables such as big bang nucleosynthesis, cosmic microwave background, and large-scale structure, the remaining objects can serve as dark matter. In addition, gravitational waves can be emitted at that time although it requires deviation from spherical symmetry in the evolution of Q-balls. Further investigation in these directions is left for future studies.